a 60kg is pushed horizontally with just enough force to start it moving across a floor and the same force continues to act afterwards. the coefficient of static friction and sliding friction 0.5 and 0.4 respectively. the acceleration of the body is :

Grade:11

1 Answers

Arun
6 years ago
See the maximum static friction that can act on the body is Umg = 0.5*60*9.8 = 294 Nnow since the force is just above 300 N, the body will translate and the type of friction now will be kinetic instead of static. Hence the net acceleration is 294-0.4*60*9.8/60 = 0.98 m/s^2 Ans.